# Affordance-Guided Reinforcement Learning via Visual Prompting

### v5

@@ Abstract
Robots equipped with reinforcement learning (RL) have the potential to learn a wide range of skills solely from a reward signal. However, obtaining a robust and dense reward signal for general manipulation tasks remains a challenge. Existing learning-based approaches require significant data, such as human demonstrations of success and failure, to learn task-specific reward functions. Recently, there is also a growing adoption of large multi-modal foundation models for robotics that can perform visual reasoning in physical contexts and generate coarse robot motions for manipulation tasks. Motivated by this range of capability, in this work, we present Keypoint-based Affordance Guidance for Improvements (KAGI), a method leveraging rewards shaped by vision-language models (VLMs) for autonomous RL. State-of-the-art VLMs have demonstrated impressive reasoning about affordances through keypoints in zero-shot, and we use these to define dense rewards that guide autonomous robotic learning. On real-world manipulation tasks specified by natural language descriptions, KAGI improves the sample efficiency of autonomous RL and enables successful task completion in 30K online fine-tuning steps. Additionally, we demonstrate the robustness of KAGI to reductions in the number of in-domain demonstrations used for pre-training, reaching similar performance in 45K online fine-tuning steps.222Project page: sites.google.com/view/affordance-guided-rl
@@ I Introduction
Recent advances in large language models (LLMs) and vision-language models (VLMs) trained on Internet-scale data show promising results in breaking down complex language instructions into task plans [12, 13, 14, 15, 16, 17], performing visual reasoning in varied contexts [18, 19, 20], and generating coarse robot motions for simple manipulation tasks [13, 21, 22, 16, 23, 20]. However, fine-tuning large pre-trained models to perform such tasks typically requires extensive human supervision, such as teleoperated demonstrations or feedback on whether the task was successfully completed. A group of prior work uses the generalization capabilities of LLMs/VLMs to solve open sets of tasks through prompt engineering and in-context learning [24, 25, 26], but these works use external motion planners and skill primitives to generate low-level actions, which require significant engineering efforts and are not robust. Another line of work prompts LLMs/VLMs to generate rewards via code to train policies or plan motions to solve the task [27, 28, 29], though existing approaches are only applied to high-level state spaces, navigation, simulated tasks, or coarse manipulation tasks. These approaches can fall short on real-world manipulation tasks that require sophisticated spatial reasoning over visual inputs.
Defining explicit rewards using pre-trained VLMs is hence an attractive approach for learning manipulation tasks. Thus far, VLMs have mainly been used to generate sparse rewards [30, 31, 6], often leading to less efficient learning. VLMs hold much richer geometric and semantic representations that we can elicit, such as reasoning about the affordances of objects and environments. In this work, we leverage insights on effective visual prompting techniques [26, 18] to generate affordance keypoints and waypoint trajectories, from which we directly derive dense rewards to improve the efficiency of online fine-tuning with RL.
@@ II-A Foundation Models for Robotics
Planning and high-level task reasoning. The rapid development of foundation models has drawn significant attention [32], demonstrating that models trained on Internet-scale data are highly adaptable to a wide range of downstream tasks, including robotics. Foundation models have enabled high-level planning and reasoning for embodied agents [16, 33, 14, 17, 34, 35, 36], generated sub-goals [37, 29, 38], and generated coarse motions for manipulation [39, 16, 20, 21, 22, 23]. These impressive results show great potential in harnessing the rich knowledge and reasoning capabilities of foundation models for tackling robotic manipulation tasks.
Despite the rapid development of LLMs for robotic control, recent works have investigated using multi-modal models to jointly provide visual and language prompts to robotic agents [12, 20, 23, 40]. These approaches recognize that converting visual observations into language descriptions and planning solely with language loses rich information from the visual modality for scene understanding. Though LLMs can provide general priors for spatial planning and reasoning, they must be properly grounded to generate accurate responses about real-world environments [41, 42, 43, 25, 44]. Our work harnesses state-of-the-art VLMs to facilitate reasoning in both language and image domains. We leverage the image domain by computing rewards in image space to bridge high-level LLM plans with low-level actions for embodied tasks.
Spatial reasoning with VLMs. Language provides a highly natural interface for providing task instructions and specifying goals. That said, robotics relies heavily on accurately perceiving and interacting with the environment. Modern VLMs, such as GPT-4o [45] and Gemini [46], demonstrate promising capabilities combining reasoning with environment perception via visual inputs. Recent approaches have explored using VLMs for generating shaping rewards via code [27, 28], reward estimation [47, 29], or preference-based learning [48]. Our work focuses on directly obtaining rewards zero-shot from VLMs by leveraging the rich geometric and semantic representations that VLMs trained on Internet-scale data possess, such that VLM-generated dense rewards can be used effectively in closed-loop autonomous RL systems. Our approach is inspired by works that use generalizable and accurate keypoint-based reasoning of VLMs for zero-shot manipulation [26, 49, 20, 18], translating high-level plans into low-level robot actions. Beyond zero-shot open-loop manipulation, our work presents a novel application of open-vocabulary visual prompting to improve policies pre-trained on diverse data via online RL fine-tuning with VLM-defined dense rewards.
@@ II-B Autonomous Reinforcement Learning
RoboFuME [6] is a recent work leveraging a reset-free pre-train fine-tune paradigm to train manipulation policies via autonomous RL. RoboFuME pre-trains offline RL policies on diverse offline data [55, 56] and in-domain demonstrations, then fine-tunes a sparse reward classifier [57] for online policy fine-tuning. However, online RL suffers when reward signals are too sparse, and fine-tuning the reward classifier requires a substantial number of in-domain demonstrations per task, which is costly but also makes the system more brittle and less robust to generalization. Our work similarly focuses on the pre-train fine-tune paradigm, specifically on improving the efficiency and robustness of online fine-tuning, enabling effective adaptation of RL policies pre-trained on diverse datasets. To do so, inspired by recent work extracting rewards from VLMs to guide zero-shot robotic manipulation [26, 31, 30] and online adaptation [6, 58], we leverage affordance representations extracted zero-shot from VLMs to tackle the dense reward shaping problem for online RL.
@@ III-B Keypoint-Based Reward Estimation via Visual Prompting
We employ an vision language model (VLM) to estimate dense rewards to facilitate fine-tuning of a pre-trained policy. First, we prompt the VLM to select appropriate grasp and target keypoints for the task, inspired by recent visual prompting techniques [26]. Then, the estimated rewards are used to generate a coarse trajectory of how the robot should complete the task. We then assign rewards to each timestep of an RL episode based on how well it follows this trajectory.
We take the first observation in the episode , consisting of a top-down image and a side view image . Our goal is to create a candidate set of grasp and target keypoints that GPT-4o can select from. We preprocess these images by (1) passing through SAM 2 [50] to get segmentations of relevant objects and sample five points from their masks, and (2) overlaying a grid and the sampled keypoints on top of to get . For the side-view image, we augment with evenly-spaced labeled horizontal lines to get , providing depth information that is not available in the top-down view. See Figures 3(a) and 3(b) for sample processed inputs. We use this density of grid lines as [26] and preliminary tests found that this was most suitable for GPT-4o to reason with. We pass and with a language instruction and metaprompt to GPT-4o, which generates block_sequence. This sequence is a list of tuples , where is a grid point chosen from representing a position in the xy-plane from top-down and is chosen from representing height in the z-axis from the side view. Dense rewards are calculated with respect to block_sequence in the next step.
For each frame in the episode, we compute the robot position in image space. We use two fitted RANSAC regressors, one for each camera view, to compute the robot position , where is from the top-down and is from the side view. Optionally, we can use pixel trackers [59] to track a specific point on the object to additionally define rewards based on object poses. Using these coordinates, we compute the nearest block to the robot position in block_sequence, (see Figure 3(c) for an example). We compute a reward is based on the L2 distance from the robot position to the block after the closest block in block_sequence, . We use the next block to encourage progression towards the goal and avoid stagnating at the current position. We transform the distance such that reward is between and : . The reward is when the sparse reward is , otherwise it is our dense reward. This reward is then optimized using an online RL algorithm [60].
@@ IV-A Experimental Setup
Task | BC | Offline RL | RoboFuME | KAGI (Ours) |
@@ IV-C Real-World Fine-Tuning with Dense Rewards
In our real-world experiments, we evaluate RoboFuME and KAGI after fine-tuning for K online environment training steps, with rewards computed in pixel image space. The results are shown in Table I. For each task, RoboFuME fine-tuning improves over the offline RL policies by -%, which is consistent with the performance in [6] and confirms that further online fine-tuning is beneficial. Notably, we see lower success rates on Cube Stacking, which requires high precision and is considerably harder than all the pick-place tasks in RoboFuME [6]. With KAGI, task performance is similar or further increases across the board.
@@ IV-D Robustness to Reduced In-Domain Demonstrations
Both RoboFuME and KAGI pre-train on a set of in-domain demonstrations for each new task, which is a nontrivial cost. We investigate the effect of reducing the number of in-domain demonstrations used by both methods. We pre-train an offline RL policy on fewer than the standard quantity of in-domain demonstrations ( forward tasks, backward tasks, and failures). We then fine-tune this policy online for K steps with RoboFuME and with KAGI. The results are shown in Table II.
Task | Offline RL | RoboFuME | KAGI (Ours) |
@@
Impact of reduced demos on the policy. Unsurprisingly, the offline RL policy performs worse than when trained on the standard quantity of in-domain demonstrations. Fine-tuning for K steps using RoboFuME somewhat improves success rate over offline RL policies, but we see a significant gap with less demos compared to its original performance with the standard number of demonstrations. Therefore, the performance of RoboFuME depends heavily on the amount of in-domain demonstrations provided. However, across all tasks, KAGI reaches close to its original performance even with this reduction. This small drop indicates that, even with fewer in-domain demonstrations, KAGI can recover comparable performance with more fine-tuning, as it uses dense shaping rewards to guide its behavior. Since KAGI can tolerate fewer task-specific demonstrations, it can be applied to new tasks more easily, with a smaller data burden.
@@ IV-E Importance of Online Fine-tuning
Since the VLM prompting component of KAGI is inspired by MOKA [26], we additionally evaluate MOKA to understand whether online fine-tuning offers any advantages. MOKA [26] leverages GPT-4V to predict affordance keypoints and plan a sequence of action primitives (e.g., ‘lift’, ‘reach grasp’, ‘grasp’), then directly executes the action primitive sequence by computing the actions required to reach each point. We evaluate two versions of MOKA: the first with precise resets that closely match the inputs to the VLM, and the second with slight perturbations to these initial conditions that is representative of environment resets after backward policy rollouts. The comparison seeks to highlight the benefits of closed-loop RL policies over open-loop primitives. In Table III, we see that MOKA with precise resets succeeds between to of the time depending on the task, which actually outperforms KAGI. However, without these precise resets, the performance of MOKA drops noticeably, as MOKA uses pre-defined motion primitives that are not robust to even slight environment perturbations, since actions are computed directly with respect to specific VLM-generated keypoints on the objects. However, KAGI can recover most of this performance loss with online fine-tuning using VLM-generated waypoint trajectories. KAGI leverages the benefits of closed-loop systems that are more generalizable, robust to environment changes, and exhibit retry behavior, as opposed to open-loop systems that use hand-designed action primitives. This makes KAGI more suitable for the autonomous RL paradigm, which involves potentially imperfect policy rollouts on a real robot.
@@ V Conclusion & Future Work
There are two key areas for future work. First, sparse task completion rewards only track if the object reached the target location or not. While this approach was successful in KAGI, incorporating object-centric waypoint trajectories into dense reward computation could further expedite learning, especially for longer-horizon tasks. Future work on improving object-centric pixel tracking and occlusion handling would aid this approach. Second, further research into improving offline RL algorithms (CalQL) could improve the performance of base policies that are fine-tuned with dense and sparse rewards, further reducing the number of online fine-tuning steps. Overall, these directions can help efficiently scale autonomous RL systems to a greater diversity of complex tasks with different objects and environments.
@@
Qualitative Analysis of VLM Performance. The performance of our online RL fine-tuning system is influenced by the accuracy of VLM keypoint and waypoint predictions. While MOKA [26] verifies that keypoint-based reasoning of VLMs are reliably accurate for tabletop manipulation and generalizable to new objects and tasks, the paper only provides a single top-down viewpoint to the VLM, since MOKA’s open-loop primitives compute actions based on pre-determined depth information specific to the evaluation setup, which requires manual engineering and tuning.
We observe that for closed-loop RL systems, providing dual-angle viewpoints are important to accurately generate waypoint trajectories in 3D space for dense rewards. We modify MOKA’s [26] metaprompt to accommodate this additional dimension of input, so as to generate 3D waypoints instead of 2D waypoints for KAGI’s dense rewards. We keep this metaprompt the same across all tasks.
To test the robustness of our prompting structure, in addition to GPT-4o, we analyze the outputs generated from our prompts when provided to GPT-4V (used by MOKA [26], larger but slower than GPT-4o) and Gemini 1.5. Qualitative analysis of each VLM’s outputs on our tasks determines that Gemini performs similarly and can replace GPT-4o in our framework. GPT-4V, while generally reliable, sometimes struggles to generate sensible depth movements using the side-angle view to facilitate successful completion of the task. Overall, KAGI is designed to flexibly incorporate existing and future VLMs of similar or better reasoning capability compared to GPT-4o, and we leave extensive qualitative comparisons and benchmarking frameworks for VLMs on 3D spatial reasoning tasks to future work.

### v6

@@ Abstract
Robots equipped with reinforcement learning (RL) have the potential to learn a wide range of skills solely from a reward signal. However, obtaining a robust and dense reward signal for general manipulation tasks remains a challenge. Existing learning-based approaches require significant data, such as human demonstrations of success and failure, to learn task-specific reward functions. Recently, there is also a growing adoption of large multi-modal foundation models for robotics that can perform visual reasoning in physical contexts and generate coarse robot motions for manipulation tasks. Motivated by this range of capability, in this work, we present Keypoint-based Affordance Guidance for Improvements (KAGI), a method leveraging rewards shaped by vision-language models (VLMs) for autonomous RL. State-of-the-art VLMs have demonstrated impressive zero-shot reasoning about affordances through keypoints, and we use these to define dense rewards that guide autonomous robotic learning. On diverse real-world manipulation tasks specified by natural language descriptions, KAGI improves the sample efficiency of autonomous RL and enables successful task completion in 30K online fine-tuning steps. Additionally, we demonstrate the robustness of KAGI to reductions in the number of in-domain demonstrations used for pre-training, reaching similar performance in 45K online fine-tuning steps.222Project page: sites.google.com/view/affordance-guided-rl
@@ I Introduction
Recent advances in large language models (LLMs) and vision-language models (VLMs) trained on Internet-scale data show promising results in breaking down complex language instructions into task plans [12, 13, 14, 15, 16, 17], performing visual reasoning in varied contexts [18, 19, 20], and generating coarse robot motions for simple manipulation tasks [13, 21, 16, 22, 20]. However, fine-tuning large pre-trained models to perform such tasks typically requires extensive human supervision, such as teleoperated demonstrations or feedback on whether the task was successfully completed.
A group of prior work uses the generalization capabilities of LLMs/VLMs to solve open sets of tasks through prompt engineering and in-context learning [23, 24], but these works use external motion planners and skill primitives to generate low-level actions, which require significant engineering efforts and are not robust. Another line of work prompts LLMs/VLMs to generate rewards via code to train policies or plan motions to solve the task [25, 26, 27]. However, existing approaches are only applied to high-level state spaces, navigation, simulated tasks, or coarse manipulation tasks, and can fall short on real-world manipulation tasks that require sophisticated spatial reasoning over visual inputs.
Defining explicit rewards using pre-trained VLMs is hence an attractive approach for learning manipulation tasks. Thus far, VLMs have mainly been used to generate sparse rewards [28, 29, 6], often leading to less efficient learning. VLMs hold much richer geometric and semantic representations that we can elicit, such as reasoning about the affordances of objects and environments. In this work, we leverage insights on effective visual prompting techniques [18] to generate affordance keypoints and waypoint trajectories, from which we directly derive dense rewards to improve the efficiency of online fine-tuning with RL.
@@ II-A Foundation Models for Robotics
Planning and high-level task reasoning. The rapid development of foundation models has drawn significant attention [30], demonstrating that models trained on Internet-scale data are highly adaptable to a wide range of downstream tasks, including robotics. Foundation models have enabled high-level planning and reasoning for embodied agents [16, 31, 14, 17, 32, 33, 34], generated sub-goals [35, 27, 36], and generated coarse motions for manipulation [37, 16, 20, 21, 22]. These impressive results show great potential in harnessing the rich knowledge and reasoning capabilities of foundation models for tackling robotic manipulation tasks.
Despite the rapid development of LLMs for robotic control, recent works have investigated using multi-modal models to jointly provide visual and language prompts to robotic agents [12, 20, 22, 38]. These approaches recognize that converting visual observations into language descriptions and planning solely with language loses rich information from the visual modality for scene understanding. Though LLMs can provide general priors for spatial planning and reasoning, they must be properly grounded to generate accurate responses about real-world environments [39, 40, 41, 24]. Our work harnesses state-of-the-art VLMs to facilitate reasoning in both language and image domains. We leverage the image domain by computing rewards in image space to bridge high-level LLM plans with low-level actions for embodied tasks.
Spatial reasoning with VLMs. Language provides a highly natural interface for providing task instructions and specifying goals. That said, robotics relies heavily on accurately perceiving and interacting with the environment. Modern VLMs, such as GPT-4o [42] and Gemini [43], demonstrate promising capabilities combining reasoning with environment perception via visual inputs. Recent approaches have explored using VLMs for shaping rewards via code [25, 26], reward estimation [44, 27], or preference-based learning [45]. Our work focuses on directly obtaining rewards zero-shot from VLMs by leveraging the rich geometric and semantic representations that VLMs trained on Internet-scale data possess, such that VLM-generated dense rewards can be used effectively in closed-loop autonomous RL systems. Our approach is inspired by works that use generalizable and accurate keypoint-based reasoning of VLMs for zero-shot manipulation [46, 20, 18], translating high-level plans into low-level robot actions. Beyond zero-shot open-loop manipulation, our work presents a novel application of open-vocabulary visual prompting, which enables us to improve policies pre-trained on diverse data via online RL fine-tuning with VLM-defined dense rewards.
@@ II-B Autonomous Reinforcement Learning
RoboFuME [6] is a recent work leveraging a reset-free pre-train fine-tune paradigm to train manipulation policies via autonomous RL. RoboFuME pre-trains offline RL policies on diverse offline data [52, 53] and in-domain demonstrations, then fine-tunes a sparse reward classifier [54] for online policy fine-tuning. However, online RL suffers when reward signals are too sparse, and fine-tuning the reward classifier requires a substantial number of in-domain demonstrations per task, which is costly but also makes the system more brittle and less robust to generalization. Our work similarly focuses on the pre-train fine-tune paradigm, specifically on improving the efficiency and robustness of online fine-tuning, enabling effective adaptation of RL policies pre-trained on diverse datasets. Inspired by recent work extracting rewards from VLMs to guide zero-shot robotic manipulation [29, 28] and online adaptation [6, 55] , we leverage affordance representations extracted zero-shot from VLMs to tackle the dense reward shaping problem for online RL.
@@ III-B Keypoint-Based Reward Estimation via Visual Prompting
We employ an vision language model (VLM) to estimate dense rewards to facilitate fine-tuning of a pre-trained policy. First, we prompt the VLM to select appropriate grasp and target keypoints for the task, inspired by recent visual prompting techniques. Then, the estimated rewards are used to generate a coarse trajectory of how the robot should complete the task. We then assign rewards to each timestep of an RL episode based on how well it follows this trajectory.
We take the first observation in the episode , consisting of a top-down image and a side view image . Our goal is to create a candidate set of grasp and target keypoints that GPT-4o can select from. We preprocess these images by (1) passing through Grounded SAM [56] to get segmentations of relevant objects and sample five points from their masks via farthest point sampling from the centroid, and (2) overlaying a grid and the sampled keypoints on top of to get . For the side-view image, we augment with evenly-spaced labeled horizontal lines to get , providing depth information that is not available in the top-down view. See Figures 3(a) and 3(b) for sample processed inputs. We use this density of grid lines as preliminary tests found that this was most suitable for GPT-4o to reason with. We pass and with a language instruction and metaprompt to GPT-4o, which generates block_sequence. This sequence is a list of tuples , where is a grid point chosen from representing a position in the xy-plane from top-down and is chosen from representing height in the z-axis from the side view. Dense rewards are calculated with respect to block_sequence in the next step.
For each frame in the episode, we compute the robot position in image space. We use two RANSAC regressors fitted on 50 manually collected keypoints, one for each camera view, to compute the robot position , where is from the top-down and is from the side view. Optionally, we can use pixel trackers [57] to track a specific point on the object to additionally define rewards based on object poses. Using these coordinates, we compute the nearest block to the robot position in block_sequence, (see Figure 3(c) for an example). We compute a reward is based on the L2 distance from the robot position to the block after the closest block in block_sequence, , to encourage progression towards the goal and avoid stagnating at the current position. We transform the distance such that reward is between and . The reward is when the sparse reward is , else it is our dense reward. This reward is optimized using an online RL algorithm [58].
@@ IV-A Experimental Setup
| Task | BC | CalQL | RoboFuME | KAGI (Ours) |
@@ IV-C Real-World Fine-Tuning with Dense Rewards
In our real-world experiments, we evaluate RoboFuME and KAGI after fine-tuning for K online environment training steps, with rewards computed in pixel image space. The results are shown in Table I. For each task, RoboFuME fine-tuning improves over CalQL policies by -%, which is consistent with the performance in [6] and confirms that further online fine-tuning is beneficial. Notably, we see lower success rates on Cube Stacking, which requires high precision and is considerably harder than all the pick-place tasks in RoboFuME [6]. With KAGI, task performance is similar or further increases across the board.
@@ IV-D Robustness to Reduced In-Domain Demonstrations
Both RoboFuME and KAGI pre-train on a set of in-domain demonstrations for each new task, which is a nontrivial cost. We investigate the effect of reducing the number of in-domain demonstrations used by both methods. We pre-train an offline RL policy via CalQL on fewer than the standard quantity of in-domain demonstrations ( forward tasks, backward tasks, and failures). We then fine-tune this policy online for K steps with RoboFuME and with KAGI. The results are shown in Table II.
| Task | CalQL | RoboFuME | KAGI (Ours) |
@@
Impact of reduced demos on the policy. Unsurprisingly, CalQL performs worse than when trained on the standard quantity of in-domain demonstrations. Fine-tuning for K steps using RoboFuME somewhat improves success rate over offline RL policies, but we see a significant gap with less demos compared to its original performance with the standard number of demonstrations. Therefore, the performance of RoboFuME depends heavily on the amount of in-domain demonstrations provided. However, across all tasks, KAGI reaches close to its original performance even with this reduction. This small drop indicates that, even with fewer in-domain demonstrations, KAGI can recover comparable performance with more fine-tuning, as it uses dense shaping rewards to guide its behavior. Since KAGI can tolerate fewer task-specific demonstrations, it can be applied to new tasks more easily, with a smaller data burden.
@@ IV-E Importance of Online Fine-tuning
Since the VLM prompting component of KAGI is inspired by MOKA [60], we additionally evaluate MOKA to understand whether online fine-tuning offers any advantages. MOKA leverages GPT-4V to predict affordance keypoints and plan a sequence of action primitives (e.g., ‘lift’, ‘reach grasp’, ‘grasp’), then directly executes the action primitive sequence by computing the actions required to reach each point. We evaluate two versions of MOKA: the first with precise resets that closely match the inputs to the VLM, and the second with slight perturbations to these initial conditions that is representative of environment resets after backward policy rollouts. The comparison seeks to highlight the benefits of closed-loop RL policies over open-loop primitives. In Table III, we see that MOKA with precise resets succeeds between to of the time depending on the task, which actually outperforms KAGI. However, without these precise resets, the performance of MOKA drops noticeably, as MOKA uses pre-defined motion primitives that are not robust to even slight environment perturbations, since actions are computed directly with respect to specific VLM-generated keypoints on the objects. However, KAGI can recover most of this performance loss with online fine-tuning using VLM-generated waypoint trajectories. KAGI leverages the benefits of closed-loop systems that are more generalizable, robust to environment changes, and exhibit retry behavior, as opposed to open-loop systems that use hand-designed action primitives. This makes KAGI more suitable for the autonomous RL paradigm, which involves potentially imperfect policy rollouts on a real robot.
@@ V Conclusion & Future Work
There are two key areas for future work. First, sparse task completion rewards only track if the object reached the target location or not. While this was successful in KAGI, incorporating object-centric waypoint trajectories into dense reward computation via pixel trackers [57] could further expedite learning, especially for longer-horizon tasks. Future work on improving object-centric pixel tracking and occlusion handling would aid this approach. Second, further research into improving offline RL algorithms (CalQL) could improve the performance of base policies that are fine-tuned with dense and sparse rewards, further reducing the number of online fine-tuning steps. Overall, these directions can help efficiently scale autonomous RL systems to more complex tasks with a greater diversity of objects and environments.
@@
Qualitative Analysis of VLM Performance. The performance of our online RL fine-tuning system is influenced by the accuracy of VLM keypoint and waypoint predictions. While MOKA verifies that keypoint-based reasoning of VLMs are reliably accurate for tabletop manipulation and generalizable to new objects and tasks, the paper only provides a single top-down viewpoint to the VLM, since MOKA’s open-loop primitives compute actions based on pre-determined depth information specific to the evaluation setup, which requires manual engineering and tuning.
We observe that for closed-loop RL systems, providing dual-angle viewpoints are important to accurately generate waypoint trajectories in 3D space for dense rewards. Furthermore, providing in-context examples help facilitate waypoint validity. We modify MOKA’s metaprompt to accommodate this additional dimension of input, so as to generate 3D waypoints instead of 2D waypoints for KAGI’s dense rewards. We keep this metaprompt the same across all tasks. To test the robustness of our prompting structure, in addition to GPT-4o, we analyze the outputs generated from our prompts when provided to GPT-4V (used by MOKA, larger but slower than GPT-4o) and Gemini 1.5. Qualitative analysis of each VLM’s outputs on our tasks determines that Gemini performs similarly and can replace GPT-4o in our framework. GPT-4V, while generally reliable, can struggle to generate sensible depth movements using the side-angle view to facilitate successful task completion. Overall, KAGI is designed to flexibly incorporate VLMs of similar or better reasoning capability compared to GPT-4o, and we leave extensive qualitative comparisons and benchmarking frameworks for VLMs on 3D spatial reasoning tasks to future work.